OpenStudy (anonymous):

2 Selected from the group at random but assume there are 5 males and 7 females. 1.) what is the probability that 2 females are selected?___________________ 2)What is the probability that 2 males are selected?_______________________

OpenStudy (anonymous):

1.\( \huge \frac{\binom{7}{2} }{\binom{12}{2}}\)

OpenStudy (anonymous):

2.\( \huge \frac{\binom{5}{2} }{\binom{12}{2}} \)

OpenStudy (anonymous):

do i multiply 7 by 2 and 12 by 2? for the 1st step

Directrix (directrix):

Of the 7 females, we select 2 and of the 5 males we select 0. C(7,2) times C(5,0) =21. Of the 12 people, there are 66 possible random selections of 2. Probability is the number of desirable outcomes (2 females) divided by the number of possible outcomes (66 groups of 2). P(2 F) = 21/ 66 = 7/22 ==> ANS Check this, CW.

Directrix (directrix):

)What is the probability that 2 males are selected. Of the 7 females, we select 0. Of the 5 males, we select 2 for our two-person group. C(7,0) times C(5,2) = 10 groups of 2-males There are C(12,2) possible groups of 2 or 66 two person groups of any gender. P(2M) = 10/66 = 5/33 ==> ANS

Directrix (directrix):

Check these, CW, and see if they are correct.

OpenStudy (anonymous):

it is good for the second not the first

Directrix (directrix):

I was looking at the 3-dice problem. I'll look at this now.

Directrix (directrix):

Do you have P(2 F) = 21/ 66 for part (a) about the two females?

OpenStudy (anonymous):

now i do:)